\begin{thebibliography}{99}
\small \setlength{\itemsep}{-0.8mm}

\bibitem{Andrews-b} G.E. Andrews, B.C. Berndt, Ramanujan's Lost Notebook, Part II,
Springer 2009.


\bibitem{Berkovich} A. Berkovich, S.O. Warnaar, Positivity preserving transformations for $q$-binomial coefficients,
 Trans. Amer. Math. Soc. 357 (2005) 2291-2351.


\bibitem{Bressoud} D.M. Bressoud, A generalization of the Rogers--Ramanujan
identities for all moduli, J. Combin. Theory Ser. A 27 (1979),
64--68.

\bibitem{BJM} K. Bringmann, C. Jennings-Shaffer, K. Mahlburg, Proofs and
reductions of various conjectured partition identities of Kanade and
Russell, J. Reine Angew. Math. 766 (2020), 109--135.


\bibitem{Cao} Z. Cao, L. Wang, Multi-sum Rogers--Ramanujan type identities, J. Math. Anal. Appl. 522 (2023), Art. 126960.

\bibitem{Cao-b}Z. Cao,  H. Rosengren, L. Wang, On some double Nahm sums of
Zagier, J. Combin. Theory Ser. A 202 (2024), Art. 105819.

\bibitem{Garrett} K. Garrett, M.E.H. Ismail, D. Stanton, Variants of the Rogers--Ramanujan identities, Adv. in Appl. Math. 23 (1999), 274--299.

\bibitem{Gasper} G. Gasper, M. Rahman, Basic Hypergeometric Series (2nd edition), Cambridge University Press, Cambridge, 2004.


\bibitem{Nancy} Nancy S.S. Gu, H. Prodinger, One-parameter generalization of Rogers--Ramanujan type identities, Adv. in Appl. Math. 45 (2010), 149--196.


\bibitem{Kanade} S. Kanade, M.C. Russell, Staircases to analytic sum-sides
for many new integer partition identities of Rogers--Ramanujan type,
Electron. J. Combin. 26 (2019), 1--6.


\bibitem{Kur} K. Kur\c{s}ung\"{o}z, Andrews--Gordon type series for Capparelli's and G\"{o}llnitz--Gordon identities,
J. Combin. Theory Ser. A 165 (2019), 117--138.


\bibitem{Ramanujan} S. Ramanujan, The Lost Notebook and Other Unpublished
Papers, Narosa, New Delhi (1988).

\bibitem{Rosengren} H. Rosengren, Proofs of some partition identities conjectured
by Kanade and Russell, Ramanujan J. 61 (2023), 295--317.


\bibitem{Schlosser} M.J. Schlosser, Bilateral identities of the Rogers--Ramanujan type, Trans. Amer.
Math. Soc. Ser. B 10 (2023), 1119--1140.

\bibitem{Slater}  L.J. Slater, Further identities of the Rogers--Ramanujan type, Proc. Lond. Math. Soc. (2) 54 (1) (1952), 147--167.

\bibitem{Uncu} A. Uncu, W. Zudilin, Reflecting (on) the modulo 9 Kanade--Russell (conjectural) identities, arXiv: 2106.02959v3.

\bibitem{Wang} L. Wang, New proofs of some double sum Rogers-Ramanujan type identities, Ramanujan J. 62 (2023), 251--272.

\bibitem{Wang-b} L. Wang, Explicit forms and proofs of Zagier's rank three examples for Nahm's
problem, Adv. Math. 450 (2024), Art. 109743.

\end{thebibliography}

\documentclass[12pt]{article}

\usepackage{amsmath, epsfig, cite}
\usepackage{amssymb}
\usepackage{amsfonts}
\usepackage{latexsym}
\usepackage{amsthm}

\newtheorem{thm}{Theorem}[section]

\newtheorem{cor}[thm]{Corollary}

\newtheorem{lem}[thm]{Lemma}

\numberwithin{equation}{section}

\renewcommand{\thefootnote}{}

\begin{document}

\begin{center}
{\large\bf Multidimensional Rogers--Ramanujan type identities\\ with
parameters
 \footnote{The corresponding author$^*$. Email addresses: weichuanan78@163.com (C. Wei), ythainmc@163.com (G. Ruan), yuanboyu81@163.com (Y. Yu).}}
\end{center}

\renewcommand{\thefootnote}{$\dagger$}

\vskip 2mm \centerline{$^1$Chuanan Wei, $^2$Guozhu Ruan, $^1$Yuanbo
Yu$^*$}
\begin{center}
{$^1$School of Biomedical Information and Engineering\\
  Hainan Medical University, Haikou 571199, China\\
  $^2$Medical Simulation Education Center\\
Hainan Medical University, Haikou 571199, China}
\end{center}


\vskip 0.7cm \noindent{\bf Abstract.} Via the contour integral
method, we establish a reduction formula from a double series to a
single series with parameters, which not only implies Uncu and
Zudilin's two results and Cao and Wang's two results, but also is
related to Berkovich and Warnaar's equation. Similarly, we also
discover some triple-sum generalizations of  Cao and Wang's
formulas. As conclusions, several multidimensional Rogers--Ramanujan
type identities with parameters or  without parameters are given.

\vskip 3mm \noindent {\it Keywords}: Rogers--Ramanujan type
identities; Jacobi's product triple identity; contour integral
method; bisection method

 \vskip 0.2cm \noindent{\it AMS
Subject Classifications:} 33D15; 11A07; 11B65

\section{Introduction}

 For any complex numbers $x$, $q$ with $|q|<1$ and nonnegative integer $n$, define the $q$-shifted factorial
 to be
 \begin{equation*}
(x;q)_{\infty}=\prod_{k=0}^{\infty}(1-xq^k)\quad\text{and}\quad
(x;q)_n=\frac{(x;q)_{\infty}}{(xq^n;q)_{\infty}}.
 \end{equation*}
For simplicity, we usually adopt the compact notation
\begin{equation*}
(x_1,x_2,\dots,x_m;q)_{n}=(x_1;q)_{n}(x_2;q)_{n}\cdots(x_m;q)_{n},
 \end{equation*}
where $m\in\mathbb{Z}^{+}$ and $n\in\mathbb{Z}^{+}\cup\{0,\infty\}.$
Following Gasper and Rahman \cite{Gasper}, define the basic
hypergeometric series as
$$
_{r+1}\phi_{r}\left[\begin{array}{c}
a_1,a_2,\ldots,a_{r+1}\\
b_1,b_2,\ldots,b_{r}
\end{array};q,\, z
\right] =\sum_{k=0}^{\infty}\frac{(a_1,a_2,\ldots, a_{r+1};q)_k}
{(q,b_1,\ldots,b_{r};q)_k}z^k.
$$

The famous Rogers--Ramanujan identities are
\begin{align}
&
\sum_{k=0}^{\infty}\frac{q^{k^2}}{(q;q)_k}=\frac{1}{(q,q^4;q^5)_{\infty}},
\label{RR-a}
\end{align}
\begin{align}
 &
\sum_{k=0}^{\infty}\frac{q^{k^2+k}}{(q;q)_k}=\frac{1}{(q^2,q^3;q^5)_{\infty}}.
\label{RR-b}
\end{align}
The bilateral generalizations of them can be seen in the  paper
\cite{Schlosser}.

In 2019,  Kanade and Russell \cite{Kanade} proposed nine conjectured
multidimensional Rogers--Ramanujan type identities related to level
2 characters of the affine Lie algebra $A_9^{(2)}$. Five of them are
confirmed by Bringmann, Jennings-Shaffer, and Mahlburg \cite{BJM}.
Rosengren \cite{Rosengren} proved all of the nine formulas by the
contour integral method.

Recently, Uncu and Zudilin \cite{Uncu} proved the following two
interesting identities:
\begin{align}
&\sum_{j,\,k\geq0}\frac{q^{j^2+2jk+2k^2}}{(q;q)_j(q^2;q^2)_k}
=\frac{(q^3;q^3)_{\infty}^2}{(q;q)_{\infty}(q^6;q^6)_{\infty}},
\label{UZ-a}\\[2mm]
&\sum_{j,\,k\geq0}\frac{q^{j^2+2jk+2k^2+j+2k}}{(q;q)_j(q^2;q^2)_k}
=\frac{(q^6;q^6)_{\infty}^2}{(q^2;q^2)_{\infty}(q^3;q^3)_{\infty}}.
\label{UZ-b}
\end{align}
Ole Warnaar has pointed that \eqref{UZ-a} and \eqref{UZ-b} are
instances of Bressoud's results from \cite{Bressoud}. Though the
contour integral method, Wang \cite{Wang} recovered \eqref{UZ-a} and
\eqref{UZ-b} and Cao and Wang \cite[Theorem 3.8]{Cao} found the
following two formulas:
\begin{align}
&\sum_{j,\,k\geq0}\frac{q^{j^2+2jk+2k^2}}{(q;q)_j(q^2;q^2)_k}(-1)^jx^{j+k}
=(qx;q^2)_{\infty},
 \label{wang-a}
\\[2mm]
&\sum_{j,\,k\geq0}\frac{q^{j^2+2jk+2k^2+k}}{(q;q)_j(q^2;q^2)_k}x^{j+2k}
=(-qx;q)_{\infty},
  \label{wang-b}
\end{align}
where $x$ is an arbitrary complex number. For more conclusions, the
reader is referred to  \cite{Andrews-b,Cao-b,
Kur,Ramanujan,Slater,Wang-b}. Inspired by the works just mentioned,
we shall establish the following theorem.

\begin{thm}\label{thm-a}
Let $x$ and $y$ be complex numbers. Then
\begin{align}
\sum_{j,\,k\geq0}\frac{q^{j^2+2jk+2k^2-j-k}}{(q;q)_j(q^2;q^2)_k}x^{j}y^{2k}
=(y;q)_{\infty}
\sum_{k=0}^{\infty}\frac{(-x/y;q)_k}{(q;q)_k(y;q)_k}q^{\binom{k}{2}}y^k.
\label{eq:wei-a}
\end{align}
\end{thm}

Choosing $(x,y)=(q,q^{\frac{1}{2}})$ in Theorem \ref{thm-a} and then
calculating the series on the right-hand side by Ramanujan's formula
(cf. \cite[Entry 5.3.2]{Andrews-b}):
\begin{align*}
 \sum_{k=0}^{\infty}\frac{(-q;q^2)_k}{(q;q)_{2k}}q^{k^2}
=\frac{(q^6;q^6)_{\infty}^2}{(q;q)_{\infty}(q^{12};q^{12})_{\infty}},
\end{align*}
we catch hold of \eqref{UZ-a}. Fixing $(x,y)=(q^2,q^{\frac{3}{2}})$
in Theorem \ref{thm-a} and then evaluating the series on the
right-hand side by Ramanujan's another formula (cf. \cite[Entry
3.4.4]{Andrews-b}):
\begin{align*}
 \sum_{k=0}^{\infty}\frac{(-q;q^2)_k}{(q;q)_{1+2k}}q^{k^2+2k}
=\frac{(q^{12};q^{12})_{\infty}(-q^6;q^6)_{\infty}}{(q;q)_{\infty}(-q^{2};q^{2})_{\infty}},
\end{align*}
we get hold of \eqref{UZ-b}.

 Taking $(x,y)\to (-qx,-q^{\frac{1}{2}}x^{\frac{1}{2}})$ in Theorem \ref{thm-a} and then computing the
series on the right-hand side by Euler's $q$-exponential formula
(cf. \cite[Appendix (II.2)]{Gasper}):
\begin{align}
 \sum_{k=0}^{\infty}\frac{q^{\binom{k}{2}}z^{k}}{(q;q)_k}=(-z;q)_{\infty},
\label{Euler-a}
\end{align}
 we arrive at \eqref{wang-a}. Letting $(x,y)\to (qx,-qx)$ in Theorem \ref{thm-a}, we are led to \eqref{wang-b}.

When $x=0$, Theorem \ref{thm-a} produces the following special case
of  Berkovich and Warnaar \cite[Equation (3.10)]{Berkovich}.

\begin{cor}\label{cor}
Let $y$ be a complex number. Then
\begin{align*}
 \sum_{k=0}^{\infty}\frac{q^{2k^2-k}y^{2k}}{(q^2;q^2)_k}
=(y;q)_{\infty}\sum_{k=0}^{\infty}\frac{q^{\binom{k}{2}}y^k}{(q;q)_k(y;q)_k}.
\end{align*}
\end{cor}

Recall a united generalization of \eqref{RR-a} and \eqref{RR-b} due
to K. Garrett, Ismail, and Stanton \cite{Garrett}:
\begin{align}
\sum_{k=0}^{\infty}\frac{q^{k^2+mk}}{(q;q)_{k}}
=\frac{(-1)^mq^{-\binom{m}{2}}E_{m-2}(q)}{(q,q^4;q^5)_{\infty}}-\frac{(-1)^mq^{-\binom{m}{2}}D_{m-2}(q)}{(q^2,q^3;q^5)_{\infty}},
\label{GST}
\end{align}
where $m$ is an integer and the Schur polynomials $D_m(q)$ and
$E_m(q)$ are defined by
\begin{align*}
&D_m(q)=D_{m-1}(q)+q^mD_{m-2}(q),\quad D_0(q)=1,\:D_1(q)=1+q,
\\[2mm]
&\qquad E_m(q)=E_{m-1}+q^mE_{m-2}(q),\quad E_0(q)=1,\:E_1(q)=1.
\end{align*}
Letting $(q,y)\to (q^2,q^{1+2m})$ in Corollary \ref{cor} and using
\eqref{GST}, we obtain the following conclusion.

\begin{cor}\label{cor-a}
Let $m$ be an integer. Then
\begin{align}
&\sum_{k=0}^{\infty}\frac{q^{k^2+2mk}}{(q^2;q^2)_{k}(q^{1+2m};q^2)_{k}}
=\frac{(-1)^mq^{2m-2m^2}E_{m-2}(q^4)}{(q^{1+2m};q^2)_{\infty}(q^4,q^{16};q^{20})_{\infty}}
\notag\\[2mm]&\qquad\qquad\qquad\qquad\qquad\qquad
-\frac{(-1)^mq^{2m-2m^2}D_{m-2}(q^4)}{(q^{1+2m};q^2)_{\infty}(q^8,q^{12};q^{20})_{\infty}}.
\label{relation-a}
\end{align}
\end{cor}
Setting $m=0$ or $1$ in Corollary \ref{cor-a}, we recover the two
known results (cf. \cite[Equations (98) and (96)]{Slater}):
\begin{align}
&
\sum_{k=0}^{\infty}\frac{q^{k^2}}{(q;q)_{2k}}=\frac{(q^{10},q^8,q^2;q^{10})_{\infty}(q^{14},q^6;q^{20})_{\infty}}{(q;q)_{\infty}},
\label{Slater-a}
\\[2mm]
&\sum_{k=0}^{\infty}\frac{q^{k^2+2k}}{(q;q)_{1+2k}}=\frac{(q^{10},q^6,q^4;q^{10})_{\infty}(q^{18},q^2;q^{20})_{\infty}}{(q;q)_{\infty}}.
\label{Slater-b}
\end{align}
It should be pointed that Gu and Prodinger \cite[Theorem 2.6]{Nancy}
gave a one-parameter generalization of \eqref{Slater-a} and
\eqref{Slater-b}, which is different from \eqref{relation-a},
several years ago.

Subsequently, we shall display the following triple-sum
generalization of \eqref{wang-a}.

\begin{thm}\label{thm-d}
Let $x$ and $y$ be complex numbers. Then
\begin{align}\label{eq:wei-d}
\sum_{j,\,k,\,\ell\,\geq0}\frac{(x;q)_j(-x)^{k+2\ell}y^{k+\ell}}{(q;q)_j(q;q)_k(q^2;q^2)_{\ell}}
q^{j+\binom{k}{2}+\binom{j+k+2\ell}{2}}
=(qx,xy;q^2)_{\infty}(-q;q)_{\infty}.
\end{align}
\end{thm}

When $x=1$, Theorem \ref{thm-d} reduces to \eqref{wang-a} thanks to
the relation (cf. \cite[P. 24]{Gasper}):
$(q,-q,-q^2;q^2)_{\infty}=1,$ which will be utilized without
indication elsewhere.

Similarly, we shall give the following two triple-sum
generalizations of \eqref{wang-b}.

\begin{thm}\label{thm-e}
Let$x$ and $y$ be complex numbers. Then
\begin{align}
&\sum_{j,\,k,\,\ell\,\geq0}\frac{(x^2y^2;q^2)_kx^jy^{2j+2\ell}}{(q;q)_j(q^2;q^2)_k(q^2;q^2)_{\ell}}
(-1)^{j+k}q^{(j+k+\ell)(j+k+\ell-1)+\ell^2+k}
\notag\\[1mm]
&\quad=\:\frac{(q;q^2)_{\infty}}{2}\big\{(xy,-y;q)_{\infty}+(-xy,y;q)_{\infty}\big\}.
\label{eq:wei-e1}
\end{align}
\end{thm}

\begin{thm}\label{thm-f}
Let $x$ and $y$ be complex numbers. Then
\begin{align}
&\sum_{j,\,k,\,\ell\,\geq0}\frac{(x^2y^2;q^2)_kx^jy^{2j+2\ell}}{(q;q)_j(q^2;q^2)_k(q^2;q^2)_{\ell}}
(-1)^{j+k}q^{(j+k+\ell)(j+k+\ell-1)+\ell^2+3k}
\notag\\[1mm]
&\quad=\:\frac{(q;q^2)_{\infty}}{2(y/q-xy)}\big\{(xy,-y/q;q)_{\infty}-(-xy,y/q;q)_{\infty}\big\}.
\label{eq:wei-f}
\end{align}
\end{thm}

When $xy=1$, Theorems \ref{thm-e} and \ref{thm-f} both become
\eqref{wang-b}. From the two theorems, we can also deduce some new
multidimensional Rogers--Ramanujan type identities.

Taking $(x,y)\to(-xq^m,q^{-m})$ in Theorem \ref{thm-e}, there is the
following formula.

\begin{cor}\label{cor-b}
Let $x$ be a complex number and let $m$ be a nonnegative integer.
Then
\begin{align}
&\sum_{j,\,k,\,\ell\,\geq0}\frac{(x^2;q^2)_k\,x^j(-1)^{k}}{(q;q)_j(q^2;q^2)_k(q^2;q^2)_{\ell}}
q^{(j+k+\ell)(j+k+\ell-1)+\ell^2+k-m(j+2\ell)}
\notag\\[1mm]
&\quad=(-q^{-m};q)_m(-x;q)_{\infty}.
 \label{eq:wei-g}
\end{align}
\end{cor}

Letting $(x,y)\to(1,x^{\frac{1}{2}})$, $(q,x^{\frac{1}{2}}/q)$ or
$(1/q,qx^{\frac{1}{2}})$ in Theorem \ref{thm-e},
 we find the following  three conclusions.

\begin{cor}\label{cor-c}
Let $x$ be a complex number. Then
\begin{align*}
&\sum_{j,\,k,\,\ell\,\geq0}\frac{(x;q^2)_k\,x^{j+\ell}(-1)^{j+k}}{(q;q)_j(q^2;q^2)_k(q^2;q^2)_{\ell}}
q^{(j+k+\ell)(j+k+\ell-1)+\ell^2+k}=\:(q,x;q^2)_{\infty},
\\[1mm]
&\sum_{j,\,k,\,\ell\,\geq0}\frac{(x;q^2)_k\,x^{j+\ell}(-1)^{j+k}}{(q;q)_j(q^2;q^2)_k(q^2;q^2)_{\ell}}
q^{(j+k+\ell)(j+k+\ell-1)+\ell^2-j+k-2\ell}=\:(q,x;q^2)_{\infty},
\end{align*}
\begin{align*}
&\sum_{j,\,k,\,\ell\,\geq0}\frac{(x;q^2)_k\,x^{j+\ell}(-1)^{j+k}}{(q;q)_j(q^2;q^2)_k(q^2;q^2)_{\ell}}
q^{(j+k+\ell)(j+k+\ell-1)+\ell^2+j+k+2\ell}=\:(q,q^2x;q^2)_{\infty}.
\end{align*}
\end{cor}

Taking $(x,y)\to(-xq^{m-1},q^{1-m})$ in Theorem \ref{thm-f}, there
holds the following formula.

\begin{cor}\label{cor-d}
Let $x$ be a complex number and let $m$ be a nonnegative integer.
Then
\begin{align}
&\sum_{j,\,k,\,\ell\,\geq0}\frac{(x^2;q^2)_k\,x^j(-1)^{k}}{(q;q)_j(q^2;q^2)_k(q^2;q^2)_{\ell}}
q^{(j+k+\ell)(j+k+\ell-1)+\ell^2+3k-(m-1)(j+2\ell)}
\notag\\
&\quad=\frac{(-q^{-m};q)_m(-x;q)_{\infty}}{q^{-m}+x}.
 \label{eq:wei-h}
\end{align}
\end{cor}

Letting $(x,y)\to(1,x^{\frac{1}{2}})$ or
$(1/q^2,q^2x^{\frac{1}{2}})$ in Theorem \ref{thm-f}, we derive the
following two conclusions.

\begin{cor}\label{cor-e}
Let $x$ be a complex number. Then
\begin{align*}
&\sum_{j,\,k,\,\ell\,\geq0}\frac{(x;q^2)_k\,x^{j+\ell}(-1)^{j+k}}{(q;q)_j(q^2;q^2)_k(q^2;q^2)_{\ell}}
q^{(j+k+\ell)(j+k+\ell-1)+\ell^2+3k}=(q^3,x;q^2)_{\infty},
\\[1mm]
&\sum_{j,\,k,\,\ell\,\geq0}\frac{(x;q^2)_k\,x^{j+\ell}(-1)^{j+k}}{(q;q)_j(q^2;q^2)_k(q^2;q^2)_{\ell}}
q^{(j+k+\ell)(j+k+\ell-1)+\ell^2+2j+3k+4\ell}=(q^3,q^2x;q^2)_{\infty}.
\end{align*}
\end{cor}

The rest of the paper is arranged as follows. In terms of the
contour integral method, we shall prove Theorems \ref{thm-a} in
Section 2. Similarly, the proof of Theorems \ref{thm-d}-\ref{thm-f}
will be displayed in Section 3. According to the bisection method,
we shall also establish several new multidimensional
Rogers--Ramanujan type identities in Section 4.

\section{Proof of Theorems \ref{thm-a}}

Let $f(z)$ represent the coming series
$$f(z)=\sum_{k=-\infty}^{\infty}a(k)z^k.$$ We employ $[z^k]f(z)$
to stand for the coefficient of $z^k$. Then it is well-known that
\begin{align*}
\oint f(z)\frac{dz}{2\pi iz}=[z^0]f(z),
\end{align*}
where the contour is positively oriented and simple closed around
the origin.

For the aim to prove Theorem \ref{thm-a}, we need the following
lemma (cf. \cite[Equation (4.10.6)]{Gasper}).

\begin{lem}\label{lem-a}
Assume that
\begin{align*}
P(z)=\frac{(a_1z,\ldots,a_Az,b_1/z,\ldots,b_B/z;q)_{\infty}}{(c_1z,\ldots,c_Az,d_1/z,\ldots,d_D/z;q)_{\infty}}
\end{align*}
has only simple poles and $|a_1\cdots a_A/c_1\cdots c_A|<1$. Then
\begin{align*}
\oint P(z)\frac{dz}{2\pi iz}
&=\frac{(b_1c_1,\ldots,b_Bc_1,a_1/c_1,\ldots,a_A/c_1;q)_{\infty}}{(q,d_1c_1,\ldots,d_Dc_1,c_2/c_1,\ldots,c_A/c_1;q)_{\infty}}
\\[1mm]
&\quad\times\sum_{k=0}^{\infty}\frac{(d_1c_1,\ldots,d_Dc_1,qc_1/a_1,\ldots,qc_1/a_A;q)_{k}}{(q,b_1c_1,\ldots,b_Bc_1,qc_1/c_2,\ldots,qc_1/c_A;q)_{k}}
\bigg(\frac{a_1\cdots a_A}{c_1\cdots c_A}\bigg)^k
\\[1mm]
&\quad+\:idem(c_1;c_2,\ldots,c_A),
\end{align*}
where the integration is over a positively oriented unit circle such
that the origin and poles of $1/(d_1/z,\ldots,d_D/z;q)_{\infty}$ lie
inside the contour and the poles of
$1/(c_1z,\ldots,c_Az;q)_{\infty}$ lie outside the contour. Here the
notation $idem(c_1;c_2,\ldots,c_A)$ after an expression means the
sum of the $A$ expressions got from the front expression by
interchanging $c_1$ with $c_k$, where $k=2,3,\ldots,c_A$.
\end{lem}

Now we begin to prove Theorem \ref{thm-a}.

\begin{proof}
Heine's transformation formulas of $_2\phi_1$ series (cf. \cite
[Appendix III.2 and III.3]{Gasper}) read
\begin{align}
_{2}\phi_{1}\left[\begin{array}{c}
a,b\\
c
\end{array};q,\, z
\right] &=\frac{(c/a,az;q)_{\infty}}{(c,z;q)_{\infty}}
\,{_{2}\phi_{1}}\left[\begin{array}{c}
abz/c,a\\
az
\end{array};q,\, \frac{c}{a}
\right]
\label{Heine-a}\\[1mm]
&=\frac{(abz/c;q)_{\infty}}{(z;q)_{\infty}}
\,{_{2}\phi_{1}}\left[\begin{array}{c}
c/a,c/b\\
c
\end{array};q,\, \frac{abz}{c}
\right]. \label{Heine-b}
\end{align}
By means of \eqref{Heine-b}, we have
\begin{align}
&{_{2}\phi_{1}}\left[\begin{array}{c}
a,aq/c\\
aq/b
\end{array};q,\, \frac{cq}{abz}
\right] =\frac{(q/z;q)_{\infty}}{(cq/abz;q)_{\infty}}
\,{_{2}\phi_{1}}\left[\begin{array}{c}
q/b,c/b\\
aq/b
\end{array};q,\, \frac{q}{z}
\right],
\label{Heine-c}\\[1mm]
&{_{2}\phi_{1}}\left[\begin{array}{c}
b,bq/c\\
bq/a
\end{array};q,\, \frac{cq}{abz}
\right] =\frac{(q/z;q)_{\infty}}{(cq/abz;q)_{\infty}}
\,{_{2}\phi_{1}}\left[\begin{array}{c}
q/a,c/a\\
bq/a
\end{array};q,\, \frac{q}{z}
\right]. \label{Heine-d}
\end{align}

Substituting \eqref{Heine-a}, \eqref{Heine-c}, and \eqref{Heine-d}
into the three-term transformation formula of $_2\phi_1$ series (cf.
\cite [Appendix III.32]{Gasper}):
\begin{align*}
_{2}\phi_{1}\left[\begin{array}{c}
a,b\\
c
\end{array};q,\, z
\right]
&=\frac{(b,c/a,az,q/az;q)_{\infty}}{(c,b/a,z,q/z;q)_{\infty}}
\,{_{2}\phi_{1}}\left[\begin{array}{c}
a,aq/c\\
aq/b
\end{array};q,\, \frac{cq}{abz}
\right]
\\[1mm]
&\quad+\frac{(a,c/b,bz,q/bz;q)_{\infty}}{(c,a/b,z,q/z;q)_{\infty}}
\,{_{2}\phi_{1}}\left[\begin{array}{c}
b,bq/c\\
bq/a
\end{array};q,\, \frac{cq}{abz}
\right],
\end{align*}
it is routine to show that
\begin{align*}
&(c/a,az;q)_{\infty}\,{_{2}\phi_{1}}\left[\begin{array}{c}
abz/c,a\\
az
\end{array};q,\, \frac{c}{a}
\right]
\notag\\[1mm]
&\:=\frac{(a,c/b,bz,q/bz;q)_{\infty}}{(a/b,cq/abz;q)_{\infty}}
\,{_{2}\phi_{1}}\left[\begin{array}{c}
q/a,c/a\\
bq/a
\end{array};q,\, \frac{q}{z}
\right]
\notag\\[1mm]
&\quad+\frac{(b,c/a,az,q/az;q)_{\infty}}{(b/a,cq/abz;q)_{\infty}}
\,{_{2}\phi_{1}}\left[\begin{array}{c}
q/b,c/b\\
aq/b
\end{array};q,\, \frac{q}{z}
\right].
\end{align*}
Take $(a,b,c,z)\to(-x/y,x/y,0,-y^2/x)$  to obtain
\begin{align}
&(y;q)_{\infty}
\sum_{k=0}^{\infty}\frac{(-x/y;q)_k}{(q;q)_k(y;q)_k}q^{\binom{k}{2}}y^k
\notag\\[1mm]
&\:=\frac{(y,x/y,q/y;q)_{\infty}}{(-1;q)_{\infty}}
 \sum_{k=0}^{\infty}\frac{(qy/x;q)_k}{(q^2;q^2)_k}\bigg(-\frac{qx}{y^2}\bigg)^k
\notag\\[1mm]
&\quad+\frac{(-y,-x/y,-q/y;q)_{\infty}}{(-1;q)_{\infty}}
 \sum_{k=0}^{\infty}\frac{(-qy/x;q)_k}{(q^2;q^2)_k}\bigg(-\frac{qx}{y^2}\bigg)^k
 \label{eq:wei-aa}.
\end{align}
Choose $(A,B,D)=(2,1,0)$ and $(a_1,a_2,b_1,c_1,c_2)=(x,q,1,y,-y)$
 in Lemma \ref{lem-a} to gain
\begin{align}
&\oint
\frac{(xz,q,qz,1/z;q)_{\infty}}{(y^2z^2;q^2)_{\infty}}\frac{dz}{2\pi
iz}
\notag\\[1mm]
&\:=\frac{(y,x/y,q/y;q)_{\infty}}{(-1;q)_{\infty}}
 \sum_{k=0}^{\infty}\frac{(qy/x;q)_k}{(q^2;q^2)_k}\bigg(-\frac{qx}{y^2}\bigg)^k
\notag\\[1mm]
&\quad+\frac{(-y,-x/y,-q/y;q)_{\infty}}{(-1;q)_{\infty}}
 \sum_{k=0}^{\infty}\frac{(-qy/x;q)_k}{(q^2;q^2)_k}\bigg(-\frac{qx}{y^2}\bigg)^k
 \label{eq:wei-bb}.
\end{align}
The combination of \eqref{eq:wei-aa} and \eqref{eq:wei-bb} engenders
\begin{align}
\oint
\frac{(xz,q,qz,1/z;q)_{\infty}}{(y^2z^2;q^2)_{\infty}}\frac{dz}{2\pi
iz}=(y;q)_{\infty}\sum_{k=0}^{\infty}\frac{(-x/y;q)_k}{(q;q)_k(y;q)_k}q^{\binom{k}{2}}y^k.
 \label{eq:wei-cc}
\end{align}

Recall Euler's another $q$-exponential formula (cf. \cite[Appendix
(II.1)]{Gasper}) and Jacobi's product triple identity (cf.
\cite[Appendix (II.28)]{Gasper}) :
\begin{align}
&\:\:\sum_{k=0}^{\infty}\frac{z^k}{(q;q)_k}=\frac{1}{(z;q)_{\infty}},
\label{Euler-b}
\\[1mm]
&\sum_{k=-\infty}^{\infty}q^{\binom{k}{2}}z^{k}=(q,-z,-q/z;q)_{\infty}.
\label{Jacobi}
\end{align}
Employing \eqref{Euler-a}, \eqref{Euler-b}, and \eqref{Jacobi}, it
is not difficult to understand that
\begin{align}
\oint
&\frac{(xz,q,qz,1/z;q)_{\infty}}{(y^2z^2;q^2)_{\infty}}\frac{dz}{2\pi
iz}
\notag\\[1mm]
&=\oint\sum_{j=0}^{\infty}\frac{q^{\binom{j}{2}}(-xz)^j}{(q;q)_j}\sum_{k=0}^{\infty}\frac{(yz)^{2k}}{(q^2;q^2)_k}
\sum_{\ell=-\infty}^{\infty}q^{\binom{\ell}{2}}(-1/z)^{\ell}\frac{dz}{2\pi
iz}
\notag\\[1mm]
&=\sum_{j,\,k\geq0}\frac{q^{j^2+2jk+2k^2-j-k}}{(q;q)_j(q^2;q^2)_k}x^{j}y^{2k}.
\label{eq:wei-dd}
\end{align}
With the help of \eqref{eq:wei-cc} and \eqref{eq:wei-dd}, we catch
hold of \eqref{eq:wei-a}.
\end{proof}

\section{ Proof of
Theorems \ref{thm-d}-\ref{thm-f}}

For proving Theorems \ref{thm-d}-\ref{thm-f}, we draw support on the
following lemma (cf. \cite[Proposition 3.2]{Rosengren}).

\begin{lem}\label{lem-b}
Let $a,b,c,t$ be complex numbers such that $|t|<1$. Then
\begin{align*}
_{2}\phi_{1}\left[\begin{array}{c}
a,b\\
c
\end{array};q,\, t
\right]=\frac{(q;q)_{\infty}}{(c,t;q)_{\infty}}\oint
\frac{(abz,cz,qz/t,t/z;q)_{\infty}}{(az,bz,cz/t;q)_{\infty}}\frac{dz}{2\pi
iz},
\end{align*}
where the integral is over a positively oriented contour separating
the origin from all poles of the integrand.
\end{lem}

Firstly, we start to prove Theorem \ref{thm-d}.

\begin{proof}
Choosing $(a,q)\to(a^2,q^2)$ in the $q$-binomial theorem:
\begin{align}
&{_{1}\phi_{0}}\left[\begin{array}{c}
a\\
-
\end{array};q,\, t
\right] =\frac{(at;q)_{\infty}}{(t;q)_{\infty}},
  \label{Bimo}
 \end{align}
it is ordinary to see that
\begin{align*}
&{_{2}\phi_{1}}\left[\begin{array}{c}
a,-a\\
-q
\end{array};q,\, t
\right] =\frac{(a^2t;q^2)_{\infty}}{(t;q^2)_{\infty}}.
 \end{align*}
Fix $(a,b,c,t)=(y^{\frac{1}{2}},-y^{\frac{1}{2}},-q,x)$ in Lemma
\ref{lem-b} and use the above identity to deduce
\begin{align}
\oint
\frac{(-yz,-qz,q,qz/x,x/z;q)_{\infty}}{(yz^2;q^2)_{\infty}(-qz/x;q)_{\infty}}\frac{dz}{2\pi
iz}=(-q;q)_{\infty}(qx,yx;q^2)_{\infty}.
 \label{eq:wei-a3}
\end{align}
Via \eqref{Euler-a}, \eqref{Euler-b}, \eqref{Jacobi}, and
\eqref{Bimo}, we have
\begin{align}
&\oint
\frac{(-yz,-qz,q,qz/x,x/z;q)_{\infty}}{(yz^2;q^2)_{\infty}(-qz/x;q)_{\infty}}\frac{dz}{2\pi
iz}\notag\\[1mm]
&\quad=\oint\sum_{j=0}^{\infty}\frac{(x;q)_j}{(q;q)_j}\Big(-\frac{qz}{x}\Big)^j\sum_{k=0}^{\infty}\frac{q^{\binom{k}{2}}(yz)^{k}}{(q;q)_k}
\notag
\end{align}
\begin{align}
&\qquad\times\sum_{\ell=0}^{\infty}\frac{(yz^2)^{\ell}}{(q^2;q^2)_{\ell}}\sum_{m=-\infty}^{\infty}(-1)^{m}q^{\binom{m}{2}}(x/z)^{m}\frac{dz}{2\pi
iz}
\notag\\[1mm]
&\quad=\sum_{j,\,k,\,\ell\,\geq0}\frac{(x;q)_j(-x)^{k+2\ell}y^{k+\ell}}{(q;q)_j(q;q)_k(q^2;q^2)_{\ell}}
q^{j+\binom{k}{2}+\binom{j+k+2\ell}{2}}. \label{eq:wei-b3}
\end{align}
The combination of \eqref{eq:wei-a3} with \eqref{eq:wei-b3} supplies
\eqref{eq:wei-d}.
\end{proof}

Secondly, we plan to prove Theorem \ref{thm-e}.

\begin{proof}
Replace $t$ by $-t$ in  \eqref{Bimo} to obtain
\begin{align}
&{_{1}\phi_{0}}\left[\begin{array}{c}
a\\
-
\end{array};q,\, -t
\right] =\frac{(-at;q)_{\infty}}{(-t;q)_{\infty}}.
  \label{Bimo-a}
 \end{align}
The sum of \eqref{Bimo} and \eqref{Bimo-a} produces
\begin{align*}
&{_{2}\phi_{1}}\left[\begin{array}{c}
a,aq\\
q
\end{array};q^2,\, t^2
\right]=\frac{1}{2}\frac{(at;q)_{\infty}}{(t;q)_{\infty}}+\frac{1}{2}\frac{(-at;q)_{\infty}}{(-t;q)_{\infty}}.
 \end{align*}
Take $(q,a,b,c,t)\to(q^2,x,qx,q,y^2)$ in Lemma \ref{lem-b} and
utilize the last equation to gain
\begin{align}
\oint
\frac{(qx^2z,qz,q^2,q^2z/y^2,y^2/z;q^2)_{\infty}}{(xz;q)_{\infty}(qz/y^2;q^2)_{\infty}}\frac{dz}{2\pi
iz}=\frac{(q;q^2)_{\infty}}{2}\big\{(xy,-y;q)_{\infty}+(-xy,y;q)_{\infty}\big\}.
 \label{eq:wei-c3}
\end{align}

Through \eqref{Euler-a}, \eqref{Euler-b}, \eqref{Jacobi}, and
\eqref{Bimo}, we arrive at
\begin{align}
&\oint
\frac{(qx^2z,qz,q^2,q^2z/y^2,y^2/z;q^2)_{\infty}}{(xz;q)_{\infty}(qz/y^2;q^2)_{\infty}}\frac{dz}{2\pi
iz}\notag\\[1mm]
&\quad=\oint\sum_{j=0}^{\infty}\frac{(xz)^j}{(q;q)_j}\sum_{k=0}^{\infty}\frac{(x^2y^2;q^2)_k}{(q^2;q^2)_k}\Big(\frac{qz}{y^2}\Big)^k
\notag\\[1mm]
&\qquad\times\sum_{\ell=0}^{\infty}\frac{(-1)^{\ell}q^{2\binom{\ell}{2}}(qz)^{\ell}}{(q^2;q^2)_{\ell}}\sum_{m=-\infty}^{\infty}(-1)^{m}q^{2\binom{m}{2}}(y^2/z)^{m}\frac{dz}{2\pi
iz}
\notag\\[1mm]
&\quad=\sum_{j,\,k,\,\ell\,\geq0}\frac{(x^2y^2;q^2)_kx^jy^{2j+2\ell}}{(q;q)_j(q^2;q^2)_k(q^2;q^2)_{\ell}}
(-1)^{j+k}q^{(j+k+\ell)(j+k+\ell-1)+\ell^2+k}. \label{eq:wei-d3}
\end{align}
Substituting \eqref{eq:wei-d3} into \eqref{eq:wei-c3}, we get hold
of \eqref{eq:wei-e1}.
\end{proof}

Thirdly, we shall prove Theorem \ref{thm-f}.

\begin{proof}
The difference of \eqref{Bimo} and \eqref{Bimo-a} can be expressed
as
\begin{align*}
&{_{2}\phi_{1}}\left[\begin{array}{c}
aq,aq^2\\
q^3
\end{array};q^2,\, t^2
\right]=\frac{1-q}{2(1-a)t}\bigg\{\frac{(at;q)_{\infty}}{(t;q)_{\infty}}-\frac{(-at;q)_{\infty}}{(-t;q)_{\infty}}\bigg\}.
 \end{align*}
Let $(q,a,b,c,t)\to(q^2, q^2x, q^3x,q^3, y^2/q^2)$ in Lemma
\ref{lem-b} and employ the upper formula to derive
\begin{align}
&\oint
\frac{(q^5x^2z,q^3z,q^2,q^4z/y^2,y^2/zq^2;q^2)_{\infty}}{(q^2xz;q)_{\infty}(q^5z/y^2;q^2)_{\infty}}\frac{dz}{2\pi
iz}
\notag\\[1mm]
&\quad=\frac{(q;q^2)_{\infty}}{2(y/q-xy)}\big\{(xy,-y/q;q)_{\infty}-(-xy,y/q;q)_{\infty}\big\}.
 \label{eq:wei-e3}
\end{align}

In terms of \eqref{Euler-a}, \eqref{Euler-b}, \eqref{Jacobi}, and
\eqref{Bimo}, there is
\begin{align}
&\oint
\frac{(q^5x^2z,q^3z,q^2,q^4z/y^2,y^2/zq^2;q^2)_{\infty}}{(q^2xz;q)_{\infty}(q^5z/y^2;q^2)_{\infty}}\frac{dz}{2\pi
iz}\notag\\[1mm]
&\quad=\oint\sum_{j=0}^{\infty}\frac{(q^2xz)^j}{(q;q)_j}\sum_{k=0}^{\infty}\frac{(x^2y^2;q^2)_k}{(q^2;q^2)_k}\Big(\frac{q^5z}{y^2}\Big)^k
\notag
\\[1mm]
&\qquad\times\sum_{\ell=0}^{\infty}\frac{(-1)^{\ell}q^{2\binom{\ell}{2}}(q^3z)^{\ell}}{(q^2;q^2)_{\ell}}\sum_{m=-\infty}^{\infty}(-1)^{m}q^{2\binom{m}{2}}(y^2/zq^2)^{m}\frac{dz}{2\pi
iz}
\notag\\[1mm]
&\quad=\sum_{j,\,k,\,\ell\,\geq0}\frac{(x^2y^2;q^2)_kx^jy^{2j+2\ell}}{(q;q)_j(q^2;q^2)_k(q^2;q^2)_{\ell}}
(-1)^{j+k}q^{(j+k+\ell)(j+k+\ell-1)+\ell^2+3k}. \label{eq:wei-f3}
\end{align}
Substituting \eqref{eq:wei-f3} into \eqref{eq:wei-e3}, we are led to
\eqref{eq:wei-f}.
\end{proof}

\section{The bisection method and
multidimensional Rogers--Ramanujan type identities}

In this section, we shall use the bisection method to establish
several new multidimensional Rogers--Ramanujan type identities.

\begin{thm}
\begin{align}
&\sum_{j,\,k\geq0}\frac{q^{4j^2+4jk+2k^2-j}}{(q;q)_{2j}(q^2;q^2)_k}
=\frac{(q^8,-q^3,-q^5;q^8)_{\infty}}{(q^2;q^2)_{\infty}},
 \label{eq:wei-a4}
\\[1mm]
&\sum_{j,\,k\geq0}\frac{q^{4j^2+4jk+2k^2+3j+2k}}{(q;q)_{1+2j}(q^2;q^2)_k}
=\frac{(q^8,-q,-q^7;q^8)_{\infty}}{(q^2;q^2)_{\infty}}.
 \label{eq:wei-b4}
\end{align}
\end{thm}

\begin{proof}
Replace $x$ by $-x$ in  \eqref{wang-b} to achieve
\begin{align}
\sum_{j,\,k\geq0}\frac{q^{j^2+2jk+2k^2+k}}{(q;q)_j(q^2;q^2)_k}(-1)^jx^{j+2k}
=(qx;q)_{\infty}.
 \label{eq:wei-c4}
\end{align}
The sum of \eqref{wang-b} and  \eqref{eq:wei-c4} creates
\begin{align}
\sum_{j,\,k\geq0}\frac{q^{4j^2+4jk+2k^2+k}}{(q;q)_{2j}(q^2;q^2)_k}x^{2j+2k}
=\frac{1}{2}\Big\{(-qx;q)_{\infty}+(qx;q)_{\infty}\Big\}.
\label{bisection-a}
\end{align}
Notice a known relation (cf. \cite[Equations (1.2a)]{Wang}):
\begin{align}
(-q;q^2)_{\infty}+(q;q^2)_{\infty}
=\frac{2}{(q^4;q^4)_{\infty}}(q^{16},-q^{6},-q^{10};q^{16})_{\infty}
 \label{WCY-a}.
\end{align}
Combing the $x=q^{-\frac{1}{2}}$ case of \eqref{bisection-a} and
\eqref{WCY-a}, we obtain \eqref{eq:wei-a4}.

The difference of \eqref{wang-b} and  \eqref{eq:wei-c4} engenders
\begin{align}
\sum_{j,\,k\geq0}\frac{q^{4j^2+4jk+2k^2+4j+3k+1}}{(q;q)_{1+2j}(q^2;q^2)_k}x^{1+2j+2k}
=\frac{1}{2}\Big\{(-qx;q)_{\infty}-(qx;q)_{\infty}\Big\}.
\label{bisection-b}
\end{align}
Notice another known relation (cf. \cite[Equations (1.2b)]{Wang}):
\begin{align}
&(-q;q^2)_{\infty}-(q;q^2)_{\infty}
=\frac{2q}{(q^4;q^4)_{\infty}}(q^{16},-q^{2},-q^{14};q^{16})_{\infty}.
 \label{WCY-b}
\end{align}
Combing the $x=q^{-\frac{1}{2}}$ case of \eqref{bisection-b} with
\eqref{WCY-b}, we discover \eqref{eq:wei-b4}.
\end{proof}

\begin{thm}Let $m$ be a nonnegative integer. Then
\begin{align}
&\sum_{j,\,k,\,\ell\,\geq0}\frac{(q;q^2)_k(-1)^{k}}{(q;q)_{2j}(q^2;q^2)_k(q^2;q^2)_{\ell}}
q^{(2j+k+\ell)(2j+k+\ell-1)+\ell^2+j+k-2m(j+\ell)}
\notag\\[1mm]
&\quad=\frac{(-q^{-m};q)_m(q^8,-q^3,-q^5;q^8)_{\infty}}{(q^2;q^2)_{\infty}},
\label{eq:wei-d4}\\[1mm]
&\sum_{j,\,k,\,\ell\,\geq0}\frac{(q;q^2)_k(-1)^{k}}{(q;q)_{1+2j}(q^2;q^2)_k(q^2;q^2)_{\ell}}
q^{(1+2j+k+\ell)(2j+k+\ell)+\ell^2+j+k-m(1+2j+2\ell)}
\notag\\[1mm]
&\quad=\frac{(-q^{-m};q)_m(q^8,-q,-q^7;q^8)_{\infty}}{(q^2;q^2)_{\infty}}.
 \label{eq:wei-e4}
\end{align}
\end{thm}

\begin{proof}
Replace $x$ by $-x$ in  \eqref{eq:wei-g} to gain
\begin{align}
&\sum_{j,\,k,\,\ell\,\geq0}\frac{(x^2;q^2)_k\,(-x)^j(-1)^{k}}{(q;q)_j(q^2;q^2)_k(q^2;q^2)_{\ell}}
q^{(j+k+\ell)(j+k+\ell-1)+\ell^2+k-m(j+2\ell)}
\notag\\
&\quad=(-q^{-m};q)_m(x;q)_{\infty}.
 \label{eq:wei-f4}
\end{align}
According to \eqref{WCY-a} and the $x=q^{\frac{1}{2}}$ case of the
sum of \eqref{eq:wei-g} and  \eqref{eq:wei-f4}, we catch hold of
\eqref{eq:wei-d4}. In accordance with \eqref{WCY-b} and the
$x=q^{\frac{1}{2}}$ case of the difference of \eqref{eq:wei-g} and
\eqref{eq:wei-f4}, we can verify \eqref{eq:wei-e4}.
\end{proof}

\begin{thm}
\begin{align}
&\sum_{j,\,k,\,\ell\,\geq0}\frac{(q^{-1};q^2)_k(-1)^{k}}{(q;q)_{2j}(q^2;q^2)_k(q^2;q^2)_{\ell}}
q^{(2j+k+\ell)(2j+k+\ell-1)+\ell^2+j+3k+2\ell}
\notag\\[1mm]
&\quad=\frac{(q^8,-q^3,-q^5;q^8)_{\infty}}{(q^2;q^2)_{\infty}},
\label{eq:wei-g4}\\[1mm]
&\sum_{j,\,k,\,\ell\,\geq0}\frac{(q^{-1};q^2)_k(-1)^{k}}{(q;q)_{1+2j}(q^2;q^2)_k(q^2;q^2)_{\ell}}
q^{(1+2j+k+\ell)(2j+k+\ell)+\ell^2+j+3k+2\ell}
\notag\\[1mm]
&\quad=\frac{(q^8,-q,-q^7;q^8)_{\infty}}{(q^2;q^2)_{\infty}}.
 \label{eq:wei-h4}
\end{align}
\end{thm}

\begin{proof}
The $m=0$ case of \eqref{eq:wei-h} can be written as
\begin{align}
\sum_{j,\,k,\,\ell\,\geq0}\frac{(x^2;q^2)_k\,x^j(-1)^{k}}{(q;q)_j(q^2;q^2)_k(q^2;q^2)_{\ell}}
q^{(j+k+\ell)(j+k+\ell-1)+\ell^2+j+3k+2\ell}=(-qx;q)_{\infty}.
 \label{eq:wei-i4}
\end{align}
Replacing $x$ by $-x$ in \eqref{eq:wei-i4}, we have
\begin{align}
\sum_{j,\,k,\,\ell\,\geq0}\frac{(x^2;q^2)_k\,(-x)^j(-1)^{k}}{(q;q)_j(q^2;q^2)_k(q^2;q^2)_{\ell}}
q^{(j+k+\ell)(j+k+\ell-1)+\ell^2+j+3k+2\ell}=(qx;q)_{\infty}.
 \label{eq:wei-j4}
\end{align}
Via \eqref{WCY-a} and the $x=q^{-\frac{1}{2}}$ case of the sum of
\eqref{eq:wei-i4} and  \eqref{eq:wei-j4}, we find \eqref{eq:wei-g4}.
Through \eqref{WCY-b} and the $x=q^{-\frac{1}{2}}$ case of the
difference of \eqref{eq:wei-i4} and \eqref{eq:wei-j4}, we can
confirm \eqref{eq:wei-h4}.
\end{proof}

\begin{thm} Let $x$ be a complex number. Then
\begin{align}
&\sum_{j,\,k,\,\ell\,\geq0}\frac{(x;q^2)_k(-1)^{k}x^{-\ell}}{(q;q)_j(q^2;q^2)_k(q^2;q^2)_{\ell}}
q^{(j+k+\ell)(j+k+\ell-1)+\ell^2+j+k+2\ell}
\notag\\[1mm]
&\quad=\frac{(-qx,-q^{3}/x;q^4)_{\infty}}{(q^2;q^4)_{\infty}},
\label{eq:wei-k4}
\\[1mm]
&\sum_{j,\,k,\,\ell\,\geq0}\frac{(x;q^2)_k(-1)^{k}x^{-\ell}}{(q;q)_j(q^2;q^2)_k(q^2;q^2)_{\ell}}
q^{(j+k+\ell)(j+k+\ell-1)+\ell^2+2j+3k+4\ell}
\notag\\[1mm]
&\quad =\frac{(-q^{3}x,-q^{5}/x;q^4)_{\infty}}{(q^2;q^4)_{\infty}}.
\label{eq:wei-l4}
\end{align}
\end{thm}

\begin{proof}
Choosing $(x,y)\to(-x/q,q/x^{\frac{1}{2}})$ in \eqref{eq:wei-e1} and
using \eqref{Jacobi}, we obtain
\begin{align*}
&\sum_{j,\,k,\,\ell\,\geq0}\frac{(x;q^2)_k(-1)^{k}x^{-\ell}}{(q;q)_j(q^2;q^2)_k(q^2;q^2)_{\ell}}
q^{(j+k+\ell)(j+k+\ell-1)+\ell^2+j+k+2\ell}
\notag\\[1mm]
&\quad=\frac{(q;q^2)_{\infty}}{2}\Big\{(-x^{\frac{1}{2}},-q/x^{\frac{1}{2}};q)_{\infty}+(x^{\frac{1}{2}},q/x^{\frac{1}{2}};q)_{\infty}\Big\}
\end{align*}
\begin{align*}
&\quad=\frac{(q;q^2)_{\infty}}{2(q;q)_{\infty}}\Big\{\sum_{k=-\infty}^{\infty}q^{\binom{k}{2}}x^{\frac{k}{2}}+\sum_{k=-\infty}^{\infty}(-1)^kq^{\binom{k}{2}}x^{\frac{k}{2}}\Big\}
\notag\\[1mm]
&\quad=\frac{1}{(q^2;q^2)_{\infty}}\sum_{k=-\infty}^{\infty}q^{\binom{2k}{2}}x^{k}\\
&\quad =\frac{(-qx,-q^{3}/x;q^4)_{\infty}}{(q^2;q^4)_{\infty}}.
\end{align*}
So we get hold of \eqref{eq:wei-k4}.

 Taking $(x,y)\to(-x/q^2,q^2/x^{\frac{1}{2}})$ in \eqref{eq:wei-f} and utilizing
\eqref{Jacobi}, we achieve
\begin{align*}
&\sum_{j,\,k,\,\ell\,\geq0}\frac{(x;q^2)_k(-1)^{k}x^{-\ell}}{(q;q)_j(q^2;q^2)_k(q^2;q^2)_{\ell}}
q^{(j+k+\ell)(j+k+\ell-1)+\ell^2+2j+3k+4\ell}
\notag\\[1mm]
&\quad=\frac{(q;q^2)_{\infty}}{2(q/x^{\frac{1}{2}}+x^{\frac{1}{2}})}\Big\{(-x^{\frac{1}{2}},-q/x^{\frac{1}{2}};q)_{\infty}-(x^{\frac{1}{2}},q/x^{\frac{1}{2}};q)_{\infty}\Big\}
\notag
\\
&\quad=\frac{(q;q^2)_{\infty}}{2(q/x^{\frac{1}{2}}+x^{\frac{1}{2}})(q;q)_{\infty}}\Big\{\sum_{k=-\infty}^{\infty}q^{\binom{k}{2}}x^{\frac{k}{2}}-\sum_{k=-\infty}^{\infty}(-1)^kq^{\binom{k}{2}}x^{\frac{k}{2}}\Big\}
\notag\\[1mm]
&\quad=\frac{1}{(q/x^{\frac{1}{2}}+x^{\frac{1}{2}})(q^2;q^2)_{\infty}}\sum_{k=-\infty}^{\infty}q^{\binom{1+2k}{2}}x^{\frac{1+2k}{2}}\\
&\quad =\frac{(-q^{3}x,-q^{5}/x;q^4)_{\infty}}{(q^2;q^4)_{\infty}}.
\end{align*}
Therefore, we complete the proof of \eqref{eq:wei-l4}.
\end{proof}

\textbf{Acknowledgments}

 The work is supported by Hainan Provincial Natural Science Foundation of China (No. 124RC511) and the National Natural Science Foundation of China (No. 12071103). We thank Ole Warnaar for some valuable comments, especially for
 pointing out that Corollary \ref{cor} is a special case of
of  Berkovich and Warnaar \cite[Equation (3.10)]{Berkovich}.


\end{document}